OpenStudy (anonymous):

What is - 3/20 as a decimal and a percent?

OpenStudy (anonymous):

0.6 and 60%

OpenStudy (anonymous):

-3/20 = -15/100 = -15% -3/20 = -15/100 = -0.15

OpenStudy (anonymous):

or no sorry i read that question wrong too XD

OpenStudy (anonymous):

xD^